OpenStudy (anonymous):

Is finding the area under a curve using the fundamental theorem of calculus and evaluating the definite integral the same thing? I really don't see a difference.

zepdrix (zepdrix):

I don't really understand what you're asking. Are we talking about the FTC Part 2?\[\large \int\limits_a^b f(x) \; dx=F(b)-F(a)\]

OpenStudy (anonymous):

Yes my hw says find area of the curve using the FTC f(x)=2sqrtx on [0,4} the next problem says evaluate the definite integrals & lists some functions. Do I do the same thing in both parts?

zepdrix (zepdrix):

Hmm that's strange. Yes it should be the same process. The FTC is really just like a... definition of the process of integration. I dunno how to word that better.. lol :3

OpenStudy (anonymous):



OpenStudy (anonymous):

Thanks zepdrix

zepdrix (zepdrix):

Whut fez? XD lol